OpenStudy (anonymous):

Ivan drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 8 hours. When Ivan drove home, there was no traffic and the trip only took 5 hours. If his average rate was 21 miles per hour faster on the trip home, how far away does Ivan live from the mountains?

OpenStudy (jazilove):

I believe it would be 105 mi

OpenStudy (victoriasushchik):

let onward speed = x mph, then distance = 7x return speed = (x+18) mph, (same) distance = 5(x+18) equating the two, 7x = 5(x+18) 2x = 90 ==> x = 45 mph so distance = 7*45 = 315 miles

OpenStudy (victoriasushchik):

its 315 miles!

OpenStudy (anonymous):

no that was wrong its 280

OpenStudy (victoriasushchik):

how did you get 280??